Grade 12Mechanics

A spy plane is being tracked by a radar at t=0, its position is reported to be (100m, 200m, 1000m). 130 seconds later its position is reported to be (2500m, 1200m, 1000m). Find a unit vector in the direction of plane velocity and the magnitude of its average velocity.

himani
8 Years agoGrade 12

1 Answer

User
ApprovedApproved Tutor Answer8 Years ago
Let us visualise this problem on the Cartesian co ordinate system.Initial position A(100m,200m,1000m) and final position B(2500,1200,1000m).
Displacement vector = (2500-100)i + (1200-200)j+ (1000-1000)k= 2400i + 1000j
Magnitude of displacement=√(2400)^2+ (1000)^2=
2600m
Therefore,magnitude of velocity= displacement /time= 2600m/130s = 20m/s.
Unit vector for velocity= velocity vector/velocity magnitude = (2400mi+ 1000mj/130s) × 1/20m/s